# A Fully Analog Implementation of Model Predictive Control with Application to Buck Converters

### v1

@@ Abstract
This paper proposes a novel approach to design analog electronic circuits that implement Model Predictive Control (MPC) policies for plants described by affine models. The combination of state-of-the-art approaches to define reduced-complexity Explicit MPC (EMPC) is employed to realize an analog circuit characterized by a limited amount of low-latency and commercially available components. The practical feasibility and effectiveness of the proposed approach are demonstrated through its application in the design of an advanced controller for DC-DC Buck converters. We formally analyze the stability of the obtained system and conduct extensive numerical simulations to demonstrate that it is capable of achieving outstanding load disturbance rejection performance, outclassing standard approaches.
@@ 1 Introduction
Model Predictive Control (MPC) is a widely employed control strategy in which, at each sampling period, the control input minimizes a given cost function, subject to input and state constraints. The MPC control framework is theoretically rich [1, 2] and particularly flexible, thanks to its inherent ability to handle constraints [3].
@@
Despite these desirable features, the practical implementation of MPC is limited. First, the typical implementation of MPC imposes a significant overhead in terms of costs and circuit area due to the need to employ expensive digital hardware, including analog-to-digital converters (ADC), digital-to-analog converters (DAC), and a computing device, making the overall system too expensive for mass production, where a limited-budget design is often a primary requirement. Second, despite the rapid advancement in the development of novel algorithms for convex optimization [4, 5], the need to solve a quadratic programming (QP) optimization problem at each iteration renders the solution infeasible for systems exhibiting fast dynamics, when fast sampling is needed.
A popular approach, enabling considerable computational savings, consists in the computation of an explicit MPC (EMPC) control law. As shown in [6], the MPC control input is a piecewise affine (PWA) function, which can be computed offline, eliminating the need for online optimization. Since the pioneering work [7], several algorithms have been proposed to efficiently compute the EMPC control law, e.g., [8, 9, 10]. Still, the EMPC approach remains critical in budget-constrained and fast-sampled applications. On the one hand, the need for digital hardware persists; on the other hand, in many practical scenarios, the PWA function is defined by a large number of regions, whose real-time evaluation may remain critical. These issues are reported in, e.g., the work [11], which applies EMPC to Buck converters.
Alternatively, fully analog circuit approaches may be considered. A notable result in this direction is [12], which constructs an analog circuit that solves linear and quadratic optimization problems, thus directly applicable to MPC implementation. However, the applications of this method are limited due to the resulting circuit complexity.
@@
The first contribution of this paper is a novel, completely general methodology to design a fully analog electronic circuit implementing MPC controllers for linear systems. We begin with an EMPC design and apply state-of-the-art approaches to obtain a PWA description of the controller characterized by a limited amount of polytopic regions. Next, we define the circuital implementation using commercially available low-latency components: comparators, resistors, operational amplifiers (OP-AMPs), and one multiplexer (MUX).
@@
The second contribution is the application of the proposed approach to the control of DC-DC Buck converters. Buck converters are devices used to regulate a high, possibly fluctuating, supply voltage to a lower and constant value. The application of Buck converters ranges from biomedical applications [13] to renewable energy generation [14, 15], across electric mobility [16] and consumer electronics. The accurate regulation of the output voltage in Buck converters is fundamental for the proper operation of the devices supplied by the converter, which otherwise may be damaged or not work properly. However, the tracking accuracy of the converter is hindered by unpredictable and sudden variations of the input supply voltage (line disturbances) and of the current drawn by the load (load disturbances) [17, 18]. To improve tracking performance, several approaches have been considered. As to line disturbance attenuation, we mention the popular feedforwarding technique [19]. In contrast, the problem of attenuating load disturbance effects is more challenging. It has been addressed by methods based on the design of circuitry to modify either the feedback signal or the pulse width of the Pulse Width Modulated (PWM) signal in response to load variations; see [20, 21, 22] for recent works in this direction. However, these solutions lack solid theoretical foundations. Other approaches aim at estimating the output current disturbance to apply a proper feedforward action. Among these, we mention the extended state observer [23], generalized proportional-integral observer [24], and the load estimator-compensator [18] approaches. These solutions are often combined with optimal controller design strategies: [23] uses linear quadratic gaussian control in conjunction with the extended state observer, and [18] uses $\mathcal{H}_{\infty}$ optimal control in conjunction with the load-estimator-compensator.
[…]
@@
To apply the proposed approach to DC-DC Buck converters, we first introduce a discrete-time mathematical model of the device and its linearization. Next, to ensure proper load disturbance rejection performance, we resort to the load estimator proposed in [18], which allows for a cheap analog implementation, and apply the proposed analog EMPC design. Thanks to the adopted regions reduction strategies, we observe a significant improvement over previous works considering EMPC design on Buck (see [11]), while ensuring a cheap circuit design. Furthermore, the proposed approach leads to a less expensive and faster implementation compared to the approach proposed in [12]. Furthermore, we provide a formal proof of global and exponential stability of the unique equilibrium point of the feedback control system, based on contraction theory.
@@
Finally, we present a thorough simulation study. We perform extensive and accurate simulation experiments to demonstrate the feasibility and effectiveness of the proposed approach. The proposed analog implementation enables high-frequency sampling at $500\text{\,}\mathrm{kHz}$, which is compatible with modern Buck design and considerably improves previous solutions, e.g.,  [26, 27, 28, 29]. We analyze the performance of the system in the presence of parametric uncertainty in the value of the Buck components through Monte Carlo simulations, considering both ceramic and electrolytic capacitor cases. Next, we perform an accurate low-level simulation utilizing the manufacturer’s models of commercially available components, demonstrating the feasibility of the proposed method even in the presence of components non-idealities, such as the finite gain-bandwidth product of the OP-AMPs and delays of comparators and MUX. Our results indicate that the controlled system enjoys outstanding load disturbance rejection performances, outperforming standard methods.
@@ 1.1 Outline
Sec. 2 reviews the standard MPC problem formulation, fixing the notation for the rest of the paper. Sec. 3 expands on the proposed design method by proposing the efficient EMPC formulation and its analog circuit implementation. In Sec. 4 we apply the proposed approach to the relevant DC-DC Buck converter control problem: we introduce the system model, detail how to estimate unmeasurable quantities, and provide a rigorous stability guarantee. Sec. 5 depicts the obtained numerical results. Finally, Sec. 6 draws conclusions.
@@ 1.2 Notation
In the following, $x=[x_{i}]_{i=1}^{N}\in\mathbb{R}^{nN}$ is the vertical concatenation of the vectors $x_{i}\in\mathbb{R}^{n}$; $I_{n}\in\mathbb{R}^{n\times n}$ is the identity matrix; $\bm{0}_{n},\bm{1}_{n}\in\mathbb{R}^{n}$ are the all-0 and all-$1$ vectors, respectively; given $A\in\mathbb{R}^{n\times n}$, $\left\|x\right\|_{A}\doteq\sqrt{x^{\top}Ax}$ is the weighted $\ell_{2}$-norm of $x\in\mathbb{R}^{n}$ and $\left\|G\right\|_{A}$ is the corresponding induced matrix norm; $\otimes$ is the Kronecker product; $e_{n}^{(m)}\in\mathbb{R}^{n}$ is the $m$-th vector of the standard Euclidean basis of $\mathbb{R}^{n}$; $E_{n}^{(m)}\in\mathbb{R}^{n\times n}$ is the matrix with $1$s on the $m$-th subdiagonal and 0s elsewhere; $\odot$, $\oplus$, and $\bar{\cdot}$ are the AND, OR, and NOT operations on Boolean (logic) variables, respectively.
@@ 2 MPC Problem Formulation
Let us consider a DT affine system $\mathcal{A}$ described by
@@
where $x_{k}\in\mathbb{R}^{n}$ is the state vector, $u_{k}\in\mathbb{R}^{n_{u}}$ is the input, $\nu_{k}\in\mathbb{R}^{n_{\nu}}$ is a vector of exogenous disturbances, $y_{k}\in\mathbb{R}^{n_{y}}$ is the output signal under control, $A\in\mathbb{R}^{n\times n}$, $B\in\mathbb{R}^{n\times n_{u}}$, $B_{\nu}\in\mathbb{R}^{n\times n_{\nu}}$, $C\in\mathbb{R}^{n_{y}\times n}$, $D\in\mathbb{R}^{n_{y}\times n_{u}}$ and $D_{\nu}\in\mathbb{R}^{n_{y}\times n_{u}}$ are matrices describing the system dynamics, while $b\in\mathbb{R}^{n}$ and $d\in\mathbb{R}^{n_{y}}$ are constant affine terms.
We define, for each $k\geq 0$, the MPC optimal control problem for the plant (1) as
[…]
@@
$H_{x}\hat{x}_{i}\leq h_{x},\quad H_{u}\hat{u}_{i}\leq h_{u},$ (2d)
$i=0,\ldots,N_{p}-1.$
Here, the decision variables are the predicted state trajectory $\hat{x}=[\hat{x}_{i}]_{i=1}^{N_{p}}\in\mathbb{R}^{nN_{p}}$, the predicted output trajectory $\hat{y}=[\hat{y}_{i}]_{i=0}^{N_{p}-1}\in\mathbb{R}^{n_{y}N_{p}}$, and the input trajectory $\hat{u}=[\hat{u}_{i}]_{i=0}^{N_{p}-1}\in\mathbb{R}^{n_{u}N_{p}}$, $N_{p}$ is the prediction horizon, $J:\mathbb{R}^{n_{y}}\times\mathbb{R}^{n_{u}}\rightarrow\mathbb{R}$ is the cost function, $(H_{x},h_{u})$ and $(H_{u},h_{u})$ represent linear inequality constraints on the state and input, respectively.
We define the MPC cost function in Eq. (2a) as
[…]
$+\sum_{i=1}^{N_{p}-1}\|\hat{u}_{i}-\hat{u}_{i-1}\|_{R_{\Delta}}^{2},$ (3)
where $(u_{r},y_{r})\in(\mathbb{R}^{n_{u}},\mathbb{R}^{n_{y}})$ are constant input and output references we aim to track, and $Q,R_{\Delta}\succeq 0$, $R\succ 0$ are symmetric weighting matrices of suitable dimensions.
@@
The triple $(u_{r},x_{r},y_{r})$ is an equilibrium of the undisturbed system (1), i.e., with $\nu_{k}=\bm{0}_{n_{\nu}}$ it holds

### v2

@@ Abstract
This paper proposes a novel approach to design analog electronic circuits that implement Model Predictive Control (MPC) policies for dynamical systems described by affine models. Effective approaches to define a reduced-complexity Explicit MPC form are combined and applied to realize an analog circuit comprising a limited set of low-latency, commercially available components. The practical feasibility and effectiveness of the proposed approach are demonstrated through its application in the design of a novel MPC-based controller for DC-DC Buck converters. We formally analyze the stability of the resulting system and conduct extensive numerical simulations to demonstrate the control system’s performance in rejecting line and load disturbances.
@@ 1 Introduction
Model predictive control (MPC) is one of the most widely adopted control techniques today, due to its flexibility to adapt to plants of diverse nature and, at the same time, to handling constraints [21]. Despite these desirable features, the practical implementation of MPC is not straightforward. First, the typical implementation of MPC imposes a significant overhead in terms of costs and circuit area due to the need to employ expensive digital hardware, including analog-to-digital converters (ADC), digital-to-analog converters (DAC), and a computing device, making the overall system too expensive for mass production, where a limited-budget design is often a primary requirement. Second, despite rapid advances in the development of novel algorithms for convex optimization [14, 16], the need to solve quadratic programming (QP) problems at each iteration renders the solution infeasible for systems exhibiting fast dynamics.
A widely adopted methodology, known for achieving significant computational efficiencies, involves the formulation of an explicit MPC (EMPC) control law. As demonstrated in [5, 6], a piecewise affine (PWA) function can be computed offline and used to define the MPC control input, obviating the need for online optimization. Although several algorithms have been proposed to efficiently compute the EMPC control law [35, 33, 7], the EMPC approach still raises criticalities in budget-constrained and fast-sampled applications. On the one hand, the need for digital hardware persists; on the other hand, when the PWA function is defined by several regions, its real-time evaluation may remain critical: the parameters defining the function may exceed the hardware storage capacity, and identifying the region containing the current parameter may be computationally expensive [8]. These issues are reported, e.g., in [42], which applies EMPC to Buck converters. A viable alternative is brought by custom hardware implementations. Notable results in this direction are [41] and [39]. [41] uses a field-programmable gate array to implement an active-set algorithm to solve the QP, but its performance remains limited by the need to perform analog-to-digital and digital-to-analog conversions. On the other hand, [39] constructs an analog circuit that solves QP, thereby making it suitable for MPC implementations. However, the applicability of this method is limited by the resulting circuit complexity.
@@
Here, we present a novel, general methodology for designing a fully analog electronic circuit that implements MPC controllers for affine systems. First, we derive an explicit MPC control law and apply state-of-the-art approaches to describe the controller in terms of a limited number of regions. Next, we define the circuit implementation using commercially available low-latency components: comparators, resistors, operational amplifiers (OP-AMPs), and one multiplexer (MUX).
@@
Furthermore, the effectiveness and feasibility of our approach are validated through the application to the control of DC-DC Buck converters. Buck converters are devices that regulate a supply voltage to a lower, constant value. Buck converters are used in various applications, including the biomedical and renewable energy fields; see, e.g., [17, 36, 34]. Accurately regulating the output voltage in Buck converters is fundamental for the proper operation of the devices supplied by the converter. However, the tracking accuracy of the converter is hindered by unpredictable and sudden variations of the input supply voltage (line disturbances) and of the current drawn by the load (load disturbances) [20, 37]. To improve tracking performance, several approaches have been considered. As to line disturbance attenuation, we mention the popular feedforwarding technique [3]. In contrast, the problem of attenuating the effects of load disturbances is more challenging. It has been addressed by methods based on the design of circuitry to modify either the feedback signal or the pulse width of the Pulse Width Modulated (PWM) signal in response to load variations; see [25, 27] for recent works in this direction. However, these solutions lack solid theoretical foundations. Other approaches aim at estimating the output current disturbance to apply a proper feedforward action. Among these, we mention the extended state observer [28], generalized proportional-integral observer [40], and the load estimator-compensator [37] approaches. These solutions are often combined with optimal controller design strategies: [28] uses linear quadratic gaussian control in conjunction with the extended state observer, and [37] uses $\mathcal{H}_{\infty}$ optimal control in conjunction with the load-estimator-compensator.
For Buck converters, the MPC controller design approach is particularly compelling due to its ability to directly handle the limitation imposed by input duty-cycle saturation between 0% and 100100%. The application of MPC to Buck converters has been explored in several contributions, e.g., [18, 32, 31, 1]. These works report successful results on practical implementations, but, in all cases, the sampling frequency must be kept small to ensure convergence of the optimization algorithm. [18, 32], and [1] report switching frequencies limited to $10\text{\,}\mathrm{kHz}$$20\text{\,}\mathrm{kHz}$, and $25\text{\,}\mathrm{kHz}$, respectively. Similar considerations apply to [31], which implements the MPC using a field-programmable gate array.
@@
To apply the proposed approach to DC-DC Buck converters, we first introduce a discrete-time (DT) mathematical model of the device and its linearization. Next, to ensure proper load-disturbance rejection performance, we use the load estimator proposed in [37], which enables a low-cost analog implementation. Finally, we apply the proposed analog EMPC design: thanks to the adopted region-reduction strategies, we observe a significant improvement over previous works on EMPC design for the Buck (see [42]), thus ensuring a low-cost circuit design and a faster implementation compared to the approach proposed in [39]. Furthermore, we provide a formal proof of the stability of the feedback control system’s equilibrium point, robustly with respect to disturbances.
@@
Our controller is validated through a comprehensive simulation study to demonstrate the feasibility and effectiveness of the proposed approach. We analyze system performance under parametric uncertainty in Buck component values using Monte Carlo simulations, considering both ceramic and electrolytic capacitor cases. Next, we perform an accurate low-level simulation utilizing the manufacturer’s models of commercially available components, demonstrating the feasibility of the proposed method even in the presence of component non-idealities. The proposed analog implementation enables high-frequency sampling at $500\text{\,}\mathrm{kHz}$, which is compatible with modern Buck designs and considerably improves upon previous solutions reported in [18, 32, 31, 1]. Our results indicate that the system exhibits outstanding disturbance-rejection performance, outperforming standard methods.
@@ 1.1 Outline
Sec. 2 reviews the standard formulation of the MPC problem. Sec. 3 expands on the proposed design method by proposing the EMPC formulation and its analog circuit implementation. In Sec. 4 we apply the proposed approach to the DC-DC Buck converter control problem: we introduce the system model, detail how to estimate unmeasurable quantities, and provide a rigorous stability guarantee. Sec. 5 depicts the obtained numerical results. Finally, Sec. 6 draws conclusions.
@@ 1.2 Notation
In the following, $x=[x_{i}]_{i=1}^{N}\in\mathbb{R}^{nN}$ is the vertical concatenation of the vectors $x_{i}\in\mathbb{R}^{n}$; $I_{n}\in\mathbb{R}^{n\times n}$ is the identity matrix; $\bm{0}_{n},\bm{1}_{n}\in\mathbb{R}^{n}$ are the null and all-$1$ vectors, respectively; $\|x\|_{M}\doteq\|M^{1/2}x\|_{2}=\sqrt{x^{\top}Mx}$ is the weighted norm of vector $x$, with weighting matrix $M$, and $\|A\|_{M}=\|M^{1/2}AM^{-1/2}\|_{2}$ is the induced weighted norm of matrix $A$; $\otimes$ is the Kronecker product; $e_{n}^{(m)}\in\mathbb{R}^{n}$ is the $m$-th vector of the standard Euclidean basis of $\mathbb{R}^{n}$; $E_{n}^{(m)}\in\mathbb{R}^{n\times n}$ is the matrix with $1$s on the $m$-th subdiagonal and 0s elsewhere; $\odot$, $\oplus$, and $\overline{\,\star\,}$ are the Boolean (logic) AND, OR, and NOT operators, respectively.
@@ 2 MPC Problem Formulation
Let us consider a discrete-time (DT) affine system, i.e.,
@@
where $x\in\mathbb{R}^{n}$, $u\in\mathbb{R}^{n_{u}}$, and $y\in\mathbb{R}^{n_{y}}$ are the state, input, and output vectors, respectively, $\nu\in\mathbb{R}^{n_{\nu}}$ is a vector of exogenous disturbances, $A\in\mathbb{R}^{n\times n}$, $B\in\mathbb{R}^{n\times n_{u}}$, $B_{\nu}\in\mathbb{R}^{n\times n_{\nu}}$, $C\in\mathbb{R}^{n_{y}\times n}$, $D\in\mathbb{R}^{n_{y}\times n_{u}}$, and $D_{\nu}\in\mathbb{R}^{n_{y}\times n_{u}}$ are matrices describing the system dynamics, $b\in\mathbb{R}^{n}$ and $d\in\mathbb{R}^{n_{y}}$ are constant affine terms.
We define, for each $k\geq 0$, the MPC optimal control problem for system (1) as:
$\min_{\hat{x},\hat{u},\hat{y}}$ $\sum_{i=0}^{N_{p}-1}\!\Big(\|\hat{y}_{i}-y_{r}\|_{Q}^{2}+\|\hat{u}_{i}-u_{r}\|_{R}^{2}\Big)+\!\!\sum_{i=1}^{N_{p}-1}\!\|\hat{u}_{i}-\hat{u}_{i-1}\|_{R_{\Delta}}^{2},$ (2a)
@@
$H_{x}\hat{x}_{i}\leq h_{x},\quad H_{u}\hat{u}_{i}\leq h_{u},~~i=0,\ldots,N_{p}-1.$ (2d)
Here, the decision variables are the predicted state trajectory $\hat{x}=[\hat{x}_{i}]_{i=1}^{N_{p}}\in\mathbb{R}^{nN_{p}}$, the predicted output trajectory $\hat{y}=[\hat{y}_{i}]_{i=0}^{N_{p}-1}\in\mathbb{R}^{n_{y}N_{p}}$, and the input sequence $\hat{u}=[\hat{u}_{i}]_{i=0}^{N_{p}-1}\in\mathbb{R}^{n_{u}N_{p}}$, where $N_{p}$ is the prediction horizon.
The cost function (2a) is composed of three terms: $\|\hat{y}_{i}-y_{r}\|_{Q}^{2}$ and $\|\hat{u}_{i}-u_{r}\|_{R}^{2}$ serve as regulation terms for the predicted output $\hat{y}_{i}$ and input $\hat{u}_{i}$ towards the constant reference output $y_{r}\in\mathbb{R}^{n_{y}}$ and input $u_{r}\in\mathbb{R}^{n_{u}}$, respectively, while $\|\hat{u}_{i}-\hat{u}_{i-1}\|_{R_{\Delta}}^{2}$ penalizes the variation in time of the inputs to obtain smoother predicted trajectories (see, e.g., [12, 13]). $Q,R_{\Delta}\succeq 0$ and $R\succ 0$ are symmetric weighting matrices of suitable dimensions. We consider the following standard assumption on the reference $(u_{r},y_{r})$:
@@
The triple $(u_{r},x_{r},y_{r})$ is an equilibrium of the undisturbed system (1), i.e., with $\nu_{k}=\bm{0}_{n_{\nu}}$ it holds that